# Probabilistic Guarantees to Explicit Constructions: Local Properties of Linear Codes

### v1

@@ 3.1 Error-Correcting Codes
For integers $N,d$, $d<N$, let $\mathcal{C}_{\textrm{out}}$ be a $N,R_{\textrm{out}},\delta_{\textrm{out}}]_{\Sigma_{\textrm{out}}}$ code and let $\mathcal{C}_{\textrm{in}}$ be a $d,R_{\textrm{in}},\delta_{\textrm{in}}]_{\Sigma_{\textrm{in}}}$ code, satisfying
$\left|{\Sigma_{\textrm{out}}}\right|=\left|\mathcal{C}_{\textrm{in}}\right|=\left|\Sigma_{\textrm{in}}\right|^{R_{\textrm{in}}\cdot d}.$ (1)
Note that Eq.˜1 allows us to construct an encoding function for $\mathcal{C}_{\textrm{in}}$ that is a bijection from $\Sigma_{\textrm{out}}$} to $\mathcal{C}_{\textrm{in}}$, denoted by $\phi\colon{\Sigma_{\textrm{out}}}\rightarrow\mathcal{C}_{\textrm{in}}$. Then the concatenated code $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$ is a $Nd,R_{\textrm{out}}\cdot R_{\textrm{in}},\delta_{\textrm{out}}\cdot\delta_{\textrm{in}}]_{\Sigma_{\textrm{in}}}$ code defined as
$\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}=\left\{v\in\Sigma_{\textrm{in}}^{([N]\times[d])}\mid\exists~c\in\mathcal{C}_{\textrm{out}},\bigl(\forall i\in[N],\phi(c[i])\in\mathcal{C}_{\textrm{in}}\bigr)\land(\phi(c[i]))_{i\in[N]}=v\right\}.$
That is, for a codeword $c\in\mathcal{C}_{\textrm{out}}$, we encode $c[i]$ for every $i\in[N]$ with the encoding map $\phi$, thus producing $(\phi(c[i]))_{i\in[N]}$. This vector is the one obtained by concatenating the codewords of $\mathcal{C}_{\textrm{in}}$ corresponding to each entry of $c$. We perform this procedure for every codeword of $\mathcal{C}_{\textrm{out}}$, and collect them in the set $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$.
It is easy to see that the rate and distance of $\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}$ is equal to $R_{\textrm{out}}\cdot R_{\textrm{in}}$ and $\delta_{\textrm{out}}\cdot\delta_{\textrm{in}}$, respectively.
@@ 3.2 Alon-Edmonds-Luby (AEL) Construction
For an integer $L$ and a matrix $A\in\Sigma_{\textrm{in}}^{([N]\times[d])\times L}$, define the matrix $A^{\mathsf{proj}}\in\Sigma_{\textrm{in}}^{([N]\times[d])\times L}$ as follows: $\forall\ell\in[N],\forall i\in[d],\forall r\in[N],\forall j\in[d]$,
@@
For $\ell\in[N]$, define the matrix $A^{\mathsf{proj}}(\ell)\in\Sigma_{\textrm{in}}^{[d]\times L}$ to be the submatrix of $A^{\mathsf{proj}}$ satisfying
@@
For an integer $n$ and a vector $h\in(\Sigma_{\textrm{in}}^{d})^{n}$, we define $h_{\mathsf{fl}}\in\Sigma_{\textrm{in}}^{([n]\times[d])}$ to be the flattened vector corresponding to $h$: that is, $h_{\mathsf{fl}}[(i,j)]=h[i](j)$ for all $i\in[n],j\in[d]$.
Note that $h_{\mathsf{fl}}$ can also be viewed as a matrix having just a single column, and therefore Definition˜3.1 applies to $h_{\mathsf{fl}}$ as well. For codes $\mathcal{C}_{\textrm{out}},\mathcal{C}_{\textrm{in}}$ where $\mathcal{C}_{\textrm{out}}$ is a $N,R_{\textrm{out}},\delta_{\textrm{out}}]_{\Sigma_{\textrm{out}}}$ code and $\mathcal{C}_{\textrm{in}}$ is a $d,R_{\textrm{in}},\delta_{\textrm{in}}]_{\Sigma_{\textrm{in}}}$ code, the code $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{out}},\mathcal{C}_{\textrm{in}},G)\subseteq\left(\Sigma_{\textrm{in}}^{d}\right)^{N}$ is defined as follows
$\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{out}},\mathcal{C}_{\textrm{in}},G):=\left\{h\in\left(\Sigma_{\textrm{in}}^{d}\right)^{N}:h^{\mathsf{proj}}_{\mathsf{fl}}\in\mathcal{C}_{\textrm{out}}\circ\mathcal{C}_{\textrm{in}}\right\}.$
@@
Let $\beta,\eta,\zeta\in[0,1]$. For infinitely many integers $N$, there is a $d=O(1/\zeta\eta^{2})$, so that the following holds. There exists a bipartite expander graph $G=(V_{L},V_{R},E)$ that can be constructed in time ${\mathrm{poly}}(N)$, with $N$ vertices on each side, degree $d$ on both sides, and with the following property: for any set $Y\subseteq V_{R}$ of right-hand vertices with $|Y|=\beta N$, we have
@@
Observation 3.5 (Explicitness of AEL Procedure).
We observe that $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{out}},\mathcal{C}_{\textrm{in}},G)\subseteq\left(\Sigma_{\textrm{in}}^{d}\right)^{N}$ can be constructed in time ${\mathrm{poly}}(N)$—that is, in time polynomial in the block length of the code—provided that each of the three components, namely, $\mathcal{C}_{\textrm{in}}$, $\mathcal{C}_{\textrm{out}}$, and $G$, can themselves be constructed in time ${\mathrm{poly}}(N)$.
@@ 4 Warm Up: Construction of List-Decodable Codes
In this section, we give a proof for Theorem˜1.8. We define the required concepts and definitions in Section˜4.1, and state the proof in Section˜4.2.
@@ 4.1 Preliminaries
For integers $n,L$, where $L$ is a constant independent of $n$, rate $R\in[0,1]$ and a sufficiently small $\varepsilon>0$, alphabet size $q\geq 2^{10L/\varepsilon}$, with probability at least $1-2^{-Ln}$, a uniformly random matrix $\mathbf{G}\in{\mathbb{F}}_{q}^{Rn\times n}$ has the following property: for every agreement hypergraph $\mathcal{H}$ on a vertex set of size $\leq L+1$ that is $(R+\varepsilon/2,n)$-weakly-partition-connected, the substituted reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}(G)$ has full column rank.
@@ 4.2 Proof for List-Decoding
Recall that the bipartite graph $G$ from ˜3.4 has $N$ vertices on each side. For the sake of simplifying the exposition, we omit floor and ceiling notation in the proof. In order to utilize the results listed in Section˜4.1, we require the inner code to be defined over a finite field ${\mathbb{F}}_{q}$, and so we take $\Sigma_{\textrm{in}}={\mathbb{F}}_{q}$. Consequently, we may interpret $\mathcal{C}_{\textrm{AEL}}\subseteq\left({\mathbb{F}}_{q}^{d}\right)^{N}$ as a code over the extension field ${\mathbb{F}}_{Q}$, where $Q=q^{d}$.
Fix a list size parameter $L$, rate $R$, and slack $\varepsilon>0$. Let $q=2^{10L/\varepsilon}$. Recall that $\delta_{\textrm{out}}$ is defined to be the distance of the outer code $\mathcal{C}_{\textrm{out}}$, and take $\eta=\varepsilon/2^{(L+3)}$ and $\zeta<\delta_{\textrm{out}}/2^{(L+1)}$. Take $d=\max(O(1/\zeta\eta^{2}),1/L)$. We take $\mathcal{C}_{\textrm{in}}\subseteq{\mathbb{F}}_{q}^{d}$ to be a ${\mathbb{F}}_{q}$-linear code having rate $R_{\textrm{in}}=R$. Let $G$ be the bipartite graph from ˜3.4 having $N$ vertices on each side, and every vertex having degree $d\geq O(1/\zeta\eta^{2})$. Let $\mathcal{C}_{\textrm{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textrm{in}}d})^{N}$ be the outer code. Note that by our definition, $\phi$ is now a map of the form $\phi\colon{\mathbb{F}}_{q}^{R_{\textrm{in}}d}\rightarrow\mathcal{C}_{\textrm{in}}$. We require $\mathcal{C}_{\textrm{out}}$ to be an ${\mathbb{F}}_{q}$-linear code and $\phi$ to be a ${\mathbb{F}}_{q}$-linear map.
@@
Let $\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G$ be as defined above. Furthermore, if $\mathcal{C}_{\textrm{in}}$ is a code satisfying the property in Corollary˜4.9, then $\mathcal{C}_{\textrm{AEL}}\subseteq\left({\mathbb{F}}_{q}^{d}\right)^{N}$ is an ${\mathbb{F}}_{q}$-linear code that is $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ average-radius list-decodable.
Since both $\mathcal{C}_{\textrm{out}}$ and the map $\phi$ are ${\mathbb{F}}_{q}$-linear, it follows that $\mathcal{C}_{\textrm{AEL}}$ is itself an ${\mathbb{F}}_{q}$-linear code.
For the sake of contradiction, assume there is a vector $y\in({\mathbb{F}}_{q}^{d})^{N}$ and pairwise distinct codewords $c_{1},\ldots,c_{L+1}\in\mathcal{C}_{\textrm{AEL}}$ such that their average Hamming distance from $y$ is less than $\frac{L}{L+1}(1-R-\varepsilon)N$. That is
@@
If the agreement hypergraph corresponding to a vector $y\in({\mathbb{F}}_{q}^{d})^{N}$ and codewords $c_{1},\ldots,c_{t}\in\mathcal{C}_{\textrm{AEL}}$ is $(R+\varepsilon,N)$-weakly-partition-connected, there exists a set $L^{*}\subseteq V_{L}$ satisfying $\left|L^{*}\right|>(1-\delta_{\textrm{out}})N$ such that for every $\ell\in L^{*}$, the agreement hypergraph corresponding to the local projections $y^{\mathsf{proj}}(\ell)\in{\mathbb{F}}_{q}^{d}$, $c_{1}^{\mathsf{proj}}(\ell),\ldots,c_{t}^{\mathsf{proj}}(\ell)\in\mathcal{C}_{\textrm{in}}$ is $(R+\frac{\varepsilon}{2},d)$-weakly-partition-connected.
@@
Because $c_{i}$ and $c_{j}$ are distinct codewords belonging to $\mathcal{C}_{\textrm{AEL}}$, we know that by construction, $|S|\geq\delta_{\textrm{out}}N$. This is because $c_{i}^{\mathsf{proj}}(\ell)\neq c_{j}^{\mathsf{proj}}(\ell)$ if and only if $\phi^{-1}(c_{i}^{\mathsf{proj}}(\ell))\neq\phi^{-1}(c_{j}^{\mathsf{proj}}(\ell))$, and this holds for exactly those vertices on which the codewords in $\mathcal{C}_{\textrm{out}}$ corresponding to $c_{i},c_{j}$ differ. Upon applying Lemma˜4.11 to the agreement hypergraph corresponding to vectors $y$ and $\left\{c_{j}\colon j\in J\right\}$, there exists at least one left vertex $\ell\in S\cap L^{*}$. Because $\ell\in S$, $c_{i}^{\mathsf{proj}}(\ell)\neq c_{j}^{\mathsf{proj}}(\ell)$, and therefore, the local codewords $\left\{c_{j}^{\mathsf{proj}}(\ell):j\in J\right\}$ are not all equal. Since $\ell\in L^{*}$, the agreement hypergraph corresponding to the local projections $y^{\mathsf{proj}}(\ell)\in{\mathbb{F}}_{q}^{d}$, $c_{1}^{\mathsf{proj}}(\ell),\ldots,c_{T}^{\mathsf{proj}}(\ell)\in\mathcal{C}_{\textrm{in}}$ is $(R+\frac{\varepsilon}{2},d)$-weakly-partition-connected. This contradicts the property of $\mathcal{C}_{\textrm{in}}$ as described in Corollary˜4.9. ∎
@@
Thus for at least $(1-\left|D_{\beta}\right|\zeta)N>(1-\delta_{\textrm{out}})N$ left vertices $\ell\in V_{L}$,
@@
Let $\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G$ be as defined in Theorem˜4.10. Furthermore, let $\mathcal{C}_{\textrm{out}}$ be a code with rate $R_{\textrm{out}}=1-\varepsilon$ and distance $\delta_{\textrm{out}}\geq\varepsilon^{3}$. Denote $Q:=q^{d}$. Then $\mathcal{C}_{\textrm{AEL}}\subseteq\left({\mathbb{F}}_{Q}\right)^{N}$ is an ${\mathbb{F}}_{q}$-linear code that is $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ average-radius list-decodable, with rate $R_{\textrm{AEL}}\geq R-\varepsilon$, where $d$ satisfies
@@
Thus, $Q=\exp(2^{O(L)}/\varepsilon^{5})$. Moreover, $\mathcal{C}_{\textrm{AEL}}$ is constructible in time ${\mathrm{poly}}(N)$.
@@
The ${\mathbb{F}}_{q}$-linearity of $\mathcal{C}_{\textrm{AEL}}$ its list-decodability parameters are proven in Theorem˜4.10. The rate is given by
$R_{\textrm{AEL}}=\frac{\log_{Q}(\left|\mathcal{C}_{\textrm{AEL}}\right|)}{N}.$
Indeed, it is easy to see that $\left|\mathcal{C}_{\textrm{AEL}}\right|=\left|\mathcal{C}_{\textrm{out}}\right|=q^{R_{\textrm{in}}R_{\textrm{out}}dN}$. Because $Q=q^{d}$, a simple calculation gives $R_{\textrm{AEL}}=R_{\textrm{in}}\cdot R_{\textrm{out}}=R\cdot(1-\varepsilon)>R-\varepsilon$. The value for $d$ is obtained by recalling that $d=\max(O(1/\zeta\eta^{2}),1/L)$, $\eta=\varepsilon/(2^{L+3})$, $\zeta=\delta_{\textrm{out}}/2^{(L+1)}$, and plugging in the value for $\delta_{\textrm{out}}$ in $\zeta$. The value for $Q$ is obtained by recalling the fact that $q=2^{10L/\varepsilon}$ from Corollary˜4.9.
We note that explicit constructions of ${\mathbb{F}}_{q}$-linear codes having rate $1-\varepsilon$ and distance $\varepsilon^{3}$ that are constructible in time ${\mathrm{poly}}(N)$ can be obtained by using Tanner codes (see Corollary 11.4.8 in [GRS23]). Moreover, the graph $G$ can be constructed in time ${\mathrm{poly}}(N)$, by ˜3.4. By Corollary˜4.9, $\mathcal{C}_{\textrm{in}}$ exists and has a block length independent of $N$, therefore it can be found through brute force in constant time. Upon invoking ˜3.5, we see that $\mathcal{C}_{\textrm{AEL}}(\mathcal{C}_{\textrm{in}},\mathcal{C}_{\textrm{out}},G)$ can be constructed in time ${\mathrm{poly}}(N)$. ∎

### v2

@@ 3.1 Error-Correcting Codes
For integers $N,d$, $d<N$, let $\mathcal{C}_{\textnormal{out}}$ be a $N,R_{\textnormal{out}},\delta_{\textnormal{out}}]_{\Sigma_{\textnormal{out}}}$ code and let $\mathcal{C}_{\textnormal{in}}$ be a $d,R_{\textnormal{in}},\delta_{\textnormal{in}}]_{\Sigma_{\textnormal{in}}}$ code, satisfying
$\left|{\Sigma_{\textnormal{out}}}\right|=\left|\mathcal{C}_{\textnormal{in}}\right|=\left|\Sigma_{\textnormal{in}}\right|^{R_{\textnormal{in}}\cdot d}.$ (1)
Note that Eq.˜1 allows us to construct an encoding function for $\mathcal{C}_{\textnormal{in}}$ that is a bijection from $\Sigma_{\textnormal{out}}$} to $\mathcal{C}_{\textnormal{in}}$, denoted by $\phi\colon{\Sigma_{\textnormal{out}}}\rightarrow\mathcal{C}_{\textnormal{in}}$. Then the concatenated code $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$ is a $Nd,R_{\textnormal{out}}\cdot R_{\textnormal{in}},\delta_{\textnormal{out}}\cdot\delta_{\textnormal{in}}]_{\Sigma_{\textnormal{in}}}$ code defined as
$\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}=\left\{v\in\Sigma_{\textnormal{in}}^{([N]\times[d])}\mid\exists~c\in\mathcal{C}_{\textnormal{out}},\bigl(\forall i\in[N],\phi(c[i])\in\mathcal{C}_{\textnormal{in}}\bigr)\land(\phi(c[i]))_{i\in[N]}=v\right\}.$
That is, for a codeword $c\in\mathcal{C}_{\textnormal{out}}$, we encode $c[i]$ for every $i\in[N]$ with the encoding map $\phi$, thus producing $(\phi(c[i]))_{i\in[N]}$. This vector is the one obtained by concatenating the codewords of $\mathcal{C}_{\textnormal{in}}$ corresponding to each entry of $c$. We perform this procedure for every codeword of $\mathcal{C}_{\textnormal{out}}$, and collect them in the set $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$.
It is easy to see that the rate and distance of $\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}$ is equal to $R_{\textnormal{out}}\cdot R_{\textnormal{in}}$ and $\delta_{\textnormal{out}}\cdot\delta_{\textnormal{in}}$, respectively.
@@ 3.2 Alon-Edmonds-Luby (AEL) Construction
For an integer $L$ and a matrix $A\in\Sigma_{\textnormal{in}}^{([N]\times[d])\times L}$, define the matrix $A^{\mathsf{proj}}\in\Sigma_{\textnormal{in}}^{([N]\times[d])\times L}$ as follows: $\forall\ell\in[N],\forall i\in[d],\forall r\in[N],\forall j\in[d]$,
@@
For $\ell\in[N]$, define the matrix $A^{\mathsf{proj}}(\ell)\in\Sigma_{\textnormal{in}}^{[d]\times L}$ to be the submatrix of $A^{\mathsf{proj}}$ satisfying
@@
For an integer $n$ and a vector $h\in(\Sigma_{\textnormal{in}}^{d})^{n}$, we define $h_{\mathsf{fl}}\in\Sigma_{\textnormal{in}}^{([n]\times[d])}$ to be the flattened vector corresponding to $h$: that is, $h_{\mathsf{fl}}[(i,j)]=h[i](j)$ for all $i\in[n],j\in[d]$.
Note that $h_{\mathsf{fl}}$ can also be viewed as a matrix having just a single column, and therefore Definition˜3.1 applies to $h_{\mathsf{fl}}$ as well. For codes $\mathcal{C}_{\textnormal{out}},\mathcal{C}_{\textnormal{in}}$ where $\mathcal{C}_{\textnormal{out}}$ is a $N,R_{\textnormal{out}},\delta_{\textnormal{out}}]_{\Sigma_{\textnormal{out}}}$ code and $\mathcal{C}_{\textnormal{in}}$ is a $d,R_{\textnormal{in}},\delta_{\textnormal{in}}]_{\Sigma_{\textnormal{in}}}$ code, the code $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{out}},\mathcal{C}_{\textnormal{in}},G)\subseteq\left(\Sigma_{\textnormal{in}}^{d}\right)^{N}$ is defined as follows
$\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{out}},\mathcal{C}_{\textnormal{in}},G):=\left\{h\in\left(\Sigma_{\textnormal{in}}^{d}\right)^{N}:h^{\mathsf{proj}}_{\mathsf{fl}}\in\mathcal{C}_{\textnormal{out}}\circ\mathcal{C}_{\textnormal{in}}\right\}.$
@@
Let $\beta,\eta,\zeta\in[0,1]$. For infinitely many integers $N$, there is a $d=O(1/\zeta\eta^{2})$, so that the following holds. There exists a bipartite $\lambda$-spectral expander graph $G=(V_{L},V_{R},E)$ that can be constructed in time ${\mathrm{poly}}(N)$, with $N$ vertices on each side, degree $d$ on both sides, and with the following property: for any set $Y\subseteq V_{R}$ of right-hand vertices with $|Y|=\beta N$, we have
@@
Observation 3.5 (Explicitness of AEL Procedure, LDPC property).
We observe that $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{out}},\mathcal{C}_{\textnormal{in}},G)\subseteq\left(\Sigma_{\textnormal{in}}^{d}\right)^{N}$ can be constructed in time ${\mathrm{poly}}(N)$—that is, in time polynomial in the block length of the code—provided that each of the three components, namely, $\mathcal{C}_{\textnormal{in}}$, $\mathcal{C}_{\textnormal{out}}$, and $G$, can themselves be constructed in time ${\mathrm{poly}}(N)$.
Moreover, $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{out}},\mathcal{C}_{\textnormal{in}},G)$ is an LDPC code if $\mathcal{C}_{\textnormal{out}}$ is an LDPC code.
@@ 4 Warm Up: Construction of List-Decodable Codes
In this section, we give a proof for Theorem˜1.11. We define the required concepts and definitions in Section˜4.1, and state the proof in Section˜4.2.
@@ 4.1 Preliminaries
For integers $n,L$, where $L$ is a constant independent of $n$, rate $R\in[0,1]$ and a sufficiently small $\varepsilon>0$, alphabet size $q\geq 2^{10L/\varepsilon}$, with probability at least $1-2^{-Ln}$, a uniformly random matrix $\mathbf{G}\in{\mathbb{F}}_{q}^{Rn\times n}$ has the following property: for every agreement hypergraph $\mathcal{H}$ on a vertex set of size $\leq L+1$ that is $(R+\varepsilon/2,n)$-weakly-partition-connected, the substituted reduced intersection matrix $\mathsf{RIM}_{\mathcal{H}}(\mathbf{G})$ has full column rank.
@@ 4.2 Proof for List-Decoding
Recall that the bipartite graph $G$ from ˜3.4 has $N$ vertices on each side. For the sake of simplifying the exposition, we omit floor and ceiling notation in the proof. In order to utilize the results listed in Section˜4.1, we require the inner code to be defined over a finite field ${\mathbb{F}}_{q}$, and so we take $\Sigma_{\textnormal{in}}={\mathbb{F}}_{q}$. Consequently, we may interpret $\mathcal{C}_{\textnormal{AEL}}\subseteq\left({\mathbb{F}}_{q}^{d}\right)^{N}$ as a code over the extension field ${\mathbb{F}}_{Q}$, where $Q=q^{d}$.
Fix a list size parameter $L$, rate $R\in[0,1]$, and slack $\varepsilon\in[0,1]$. Let $q=2^{10L/\varepsilon}$. Recall that $\delta_{\textnormal{out}}$ is defined to be the distance of the outer code $\mathcal{C}_{\textnormal{out}}$, and take $\eta=\varepsilon/2^{(L+3)}$ and $\zeta<\delta_{\textnormal{out}}/2^{(L+1)}$. Take $d=\max(O(1/\zeta\eta^{2}),1/L)$. We take $\mathcal{C}_{\textnormal{in}}\subseteq{\mathbb{F}}_{q}^{d}$ to be a ${\mathbb{F}}_{q}$-linear code having rate $R_{\textnormal{in}}=R$. Let $G$ be the bipartite graph from ˜3.4 having $N$ vertices on each side, and every vertex having degree $d\geq O(1/\zeta\eta^{2})$. Let $\mathcal{C}_{\textnormal{out}}\subseteq({\mathbb{F}}_{q}^{R_{\textnormal{in}}d})^{N}$ be the outer code. Note that by our definition, $\phi$ is now a map of the form $\phi\colon{\mathbb{F}}_{q}^{R_{\textnormal{in}}d}\rightarrow\mathcal{C}_{\textnormal{in}}$. We require $\mathcal{C}_{\textnormal{out}}$ to be an ${\mathbb{F}}_{q}$-linear code and $\phi$ to be a ${\mathbb{F}}_{q}$-linear map.
@@
Let $\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G$ be as defined above. Furthermore, if $\mathcal{C}_{\textnormal{in}}$ is a code satisfying the property in Corollary˜4.9, then $\mathcal{C}_{\textnormal{AEL}}\subseteq\left({\mathbb{F}}_{q}^{d}\right)^{N}$ is an ${\mathbb{F}}_{q}$-linear code that is $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ average-radius list-decodable.
Since both $\mathcal{C}_{\textnormal{out}}$ and the map $\phi$ are ${\mathbb{F}}_{q}$-linear, it follows that $\mathcal{C}_{\textnormal{AEL}}$ is itself an ${\mathbb{F}}_{q}$-linear code.
For the sake of contradiction, assume there is a vector $y\in({\mathbb{F}}_{q}^{d})^{N}$ and pairwise distinct codewords $c_{1},\ldots,c_{L+1}\in\mathcal{C}_{\textnormal{AEL}}$ such that their average Hamming distance from $y$ is less than $\frac{L}{L+1}(1-R-\varepsilon)N$. That is
@@
If the agreement hypergraph corresponding to a vector $y\in({\mathbb{F}}_{q}^{d})^{N}$ and codewords $c_{1},\ldots,c_{t}\in\mathcal{C}_{\textnormal{AEL}}$ is $(R+\varepsilon,N)$-weakly-partition-connected, there exists a set $L^{*}\subseteq V_{L}$ satisfying $\left|L^{*}\right|>(1-\delta_{\textnormal{out}})N$ such that for every $\ell\in L^{*}$, the agreement hypergraph corresponding to the local projections $y^{\mathsf{proj}}(\ell)\in{\mathbb{F}}_{q}^{d}$, $c_{1}^{\mathsf{proj}}(\ell),\ldots,c_{t}^{\mathsf{proj}}(\ell)\in\mathcal{C}_{\textnormal{in}}$ is $(R+\frac{\varepsilon}{2},d)$-weakly-partition-connected.
@@
Because $c_{i}$ and $c_{j}$ are distinct codewords belonging to $\mathcal{C}_{\textnormal{AEL}}$, we know that by construction, $|S|\geq\delta_{\textnormal{out}}N$. This is because $c_{i}^{\mathsf{proj}}(\ell)\neq c_{j}^{\mathsf{proj}}(\ell)$ if and only if $\phi^{-1}(c_{i}^{\mathsf{proj}}(\ell))\neq\phi^{-1}(c_{j}^{\mathsf{proj}}(\ell))$, and this holds for exactly those vertices on which the codewords in $\mathcal{C}_{\textnormal{out}}$ corresponding to $c_{i},c_{j}$ differ. Upon applying Lemma˜4.11 to the agreement hypergraph corresponding to vectors $y$ and $\left\{c_{j}\colon j\in J\right\}$, there exists at least one left vertex $\ell\in S\cap L^{*}$. Because $\ell\in S$, $c_{i}^{\mathsf{proj}}(\ell)\neq c_{j}^{\mathsf{proj}}(\ell)$, and therefore, the local codewords $\left\{c_{j}^{\mathsf{proj}}(\ell):j\in J\right\}$ are not all equal. Since $\ell\in L^{*}$, the agreement hypergraph corresponding to the local projections $y^{\mathsf{proj}}(\ell)\in{\mathbb{F}}_{q}^{d}$, $c_{1}^{\mathsf{proj}}(\ell),\ldots,c_{T}^{\mathsf{proj}}(\ell)\in\mathcal{C}_{\textnormal{in}}$ is $(R+\frac{\varepsilon}{2},d)$-weakly-partition-connected. This contradicts the property of $\mathcal{C}_{\textnormal{in}}$ as described in Corollary˜4.9. ∎
@@
Thus for at least $(1-\left|D_{\beta}\right|\zeta)N>(1-\delta_{\textnormal{out}})N$ left vertices $\ell\in V_{L}$,
@@
Let $\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G$ be as defined in Theorem˜4.10. Furthermore, let $\mathcal{C}_{\textnormal{out}}$ be a code with rate $R_{\textnormal{out}}=1-\varepsilon$ and distance $\delta_{\textnormal{out}}\geq\varepsilon^{3}$. Denote $Q:=q^{d}$. Then $\mathcal{C}_{\textnormal{AEL}}\subseteq\left({\mathbb{F}}_{Q}\right)^{N}$ is an ${\mathbb{F}}_{q}$-linear LDPC code that is $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ average-radius list-decodable, with rate $R_{\textnormal{AEL}}\geq R-\varepsilon$, where $d$ satisfies
@@
Thus, $Q=\exp(2^{O(L)}/\varepsilon^{5})$. Moreover, $\mathcal{C}_{\textnormal{AEL}}$ is constructible in time ${\mathrm{poly}}(N)$.
@@
The ${\mathbb{F}}_{q}$-linearity of $\mathcal{C}_{\textnormal{AEL}}$ its list-decodability parameters are proven in Theorem˜4.10. The rate is given by
$R_{\textnormal{AEL}}=\frac{\log_{Q}(\left|\mathcal{C}_{\textnormal{AEL}}\right|)}{N}.$
Indeed, it is easy to see that $\left|\mathcal{C}_{\textnormal{AEL}}\right|=\left|\mathcal{C}_{\textnormal{out}}\right|=q^{R_{\textnormal{in}}R_{\textnormal{out}}dN}$. Because $Q=q^{d}$, a simple calculation gives $R_{\textnormal{AEL}}=R_{\textnormal{in}}\cdot R_{\textnormal{out}}=R\cdot(1-\varepsilon)>R-\varepsilon$. The value for $d$ is obtained by recalling that $d=\max(O(1/\zeta\eta^{2}),1/L)$, $\eta=\varepsilon/(2^{L+3})$, $\zeta=\delta_{\textnormal{out}}/2^{(L+1)}$, and plugging in the value for $\delta_{\textnormal{out}}$ in $\zeta$. The value for $Q$ is obtained by recalling the fact that $q=2^{10L/\varepsilon}$ from Corollary˜4.9.
We note that explicit constructions of ${\mathbb{F}}_{q}$-linear codes having rate $1-\varepsilon$ and distance $\varepsilon^{3}$ that are constructible in time ${\mathrm{poly}}(N)$ can be obtained by using Tanner codes (see Corollary 11.4.8 in [GRS23]). We note that Tanner codes are LDPC codes. Moreover, the graph $G$ can be constructed in time ${\mathrm{poly}}(N)$, by ˜3.4. By Corollary˜4.9, $\mathcal{C}_{\textnormal{in}}$ exists and has a block length independent of $N$, therefore it can be found through brute force in constant time. Upon invoking ˜3.5, we see that $\mathcal{C}_{\textnormal{AEL}}(\mathcal{C}_{\textnormal{in}},\mathcal{C}_{\textnormal{out}},G)$ is an LDPC code and can be constructed in time ${\mathrm{poly}}(N)$. ∎